# New simple and accurate approximations for quintessence

### v2

@@ 4 Numerical comparisons
The relative errors of $w_{\phi}(z)$, $\Omega_{\mathrm{m}}(z)$ and $E(z)$ are depicted in figure 1 for our $[0/1]_{w_{\mathrm{DE}}}(T)$ and $[1/1]_{w_{\mathrm{DE}}}(T)$ based approximations, as well as the approximations by Scherrer and Sen [52] and Chiba [53], given by the equations (70) and (76), respectively, in appendix A. As can be seen, the previous approximations and the $[0/1]_{w_{\mathrm{DE}}}(T)$ approximant based approximations have similar relative errors for the physical observables $\Omega_{\mathrm{m}}$ and $E_{\phi}$, with respective $\sim 3\%$ and $\sim 1.5\%$ maxima, while the $[1/1]_{w_{\mathrm{DE}}}(T)$ approximant yields a maximum relative error of $\sim 0.1\%$ for $\Omega_{\mathrm{m}}$ and an even smaller error for $E_{\phi}$, making this approximation observationally indistinguishable from numerical results. Note that changing $\phi_{*}$ so that $\lambda_{*}$ decreases also decreases the errors, which is due to slower rolling of the scalar field. Note also that the errors of all approximations decrease as $z$ increases, which should not come as a surprise since decreasing $z$ entails increased matter domination. Investigation of several other scalar field potentials yield similar results.
@@
Figure 1: Thawing quintessence: Relative errors for $w_{\phi}$, $\Omega_{\mathrm{m}}$, $E_{\phi}$. The dotted curves depict the relative errors of the new $[0/1]_{w_{\mathrm{DE}}}(T)$ based approximations (59) while the thick dotted curves represent the relative errors of the $[1/1]_{w_{\mathrm{DE}}}(T)$ based approximations (60). In Figure (a) the dashed curve is the approximation by Scherrer and Sen [52] given in eq. (70), while the space-dashed curve is the approximation obtained by Chiba [53], see eq. (76). In Fig. (b) and (c) the dashed curves correspond to the $\Lambda$CDM solution given in (19), which is the basis for the approximations of both Scherrer and Sen [52] and Chiba [53].
Tracking quintessence
Thus the derivatives of $\Gamma$ are zero and hence $\Gamma^{(1)}=0$ and $\Gamma^{(2)}=0$ in (57), while, as for all the present tracking quintessence models, $w_{\infty}=-2/(2+p)$. Thus, $w_{\infty}$, $\gamma$ and $\beta$ only involve a single independent parameter, $p$, or, equivalently, $w_{\infty}$, which yields a slightly simpler expressions for $\gamma$ and $\beta$ than $p$. Using the system (52) and the value $\Omega_{\phi 0}=0.68$, which form a surface in the $(\tilde{\phi},u,v)$ state space that intersects the tracker orbit at a point found numerically, and where the numerical calculations also give us $w_{\phi}=u^{2}-1$, $\Omega_{\mathrm{m}}=1-\Omega_{\phi}=1-3\tilde{\phi}v^{2}$, while $E_{\phi}$ is found numerically by integrating [^16] The present model, like all scalar field models, also admits a 1-parameter set of thawing quintessence solutions in the state space, see [32], which we for brevity do not discuss.
[…]
(a) Relative errors of $w_{\phi}$
(b) Relative errors of $\Omega_{\mathrm{m}}$
(c) Relative errors of $E_{\phi}$
Figure 2: Tracking quintessence: Relative errors for $w_{\phi}$, $\Omega_{\mathrm{m}}$, $E_{\phi}$. The dotted curves depict the relative errors for the $[0/1]_{w_{\mathrm{DE}}}(T)$ based approximations given by (59), while the thick dotted curves represent the relative errors for the $[1/1]_{w_{\mathrm{DE}}}(T)$ based approximations (60). In Figure (a) the space-dashed and dashed curves describe the Chiba approximation [54], given in (79), truncated at 1st and 2nd order in $\Omega_{\phi}$, respectively. In Fig. (b) and (c) the dashed curves depict the $w$CDM solution given in (19), which form the basis for the Chiba approximations.
@@ 5 Discussion
In this paper we have introduced simple and accurate approximations for thawing and tracking quintessence situated in a broader DE context. This was accomplished by identifying that thawing and tracking quintessence are associated with ‘attractor solutions’ for which $\lim_{N\rightarrow-\infty}\Omega_{\phi}=0$ (and hence $\lim_{N\rightarrow-\infty}\Omega_{\mathrm{m}}=1$) and $\lim_{N\rightarrow-\infty}w_{\phi}=w_{\infty}$, where $w_{\infty}=-1$ for thawing quintessence while $-1<w_{\infty}<0$ for tracking quintessence. These asymptotic properties served as motivation for situating the results in a more general DE setting, which resulted in series expansions in $T=T_{0}\exp(-3w_{\infty}N)$. To improve the accuracy we used Padé approximants where the $[1/1]$ Padé approximant for $\Omega_{\mathrm{DE}}$ was used to relate the approximations to the present time initial value $\Omega_{\mathrm{DE}0}$, or, equivalently, $\Omega_{\mathrm{m}0}=\rho_{\mathrm{m}0}/3H_{0}^{2}$.
We note that the $[0/1]_{w_{\mathrm{DE}}}(T)$ Padé approximant contains as many model parameters ($w_{\infty}$ and $\gamma$) as CPL ($w_{0}$, or equivalently $w_{\infty}$, and $w_{a}$) while the $[1/1]_{w_{\mathrm{DE}}}(T)$ approximant has an additional model parameter $\beta$. However, some potentials yield parameter relationships (or one can impose relationships in the DE context) between $w_{\infty}$, $\gamma$ and $\beta$; notably the inverse power law potential, $V\propto\phi^{-p}$, yields $w_{\infty}(p)$, $\gamma(p)$, $\beta(p)$ and thus our approximations for these models contain one less model parameter than CPL; note, however, that to obtain a specific spacetime, one, in addition to model parameters, needs to specify the initial data $H_{0}$ and $\rho_{\mathrm{m}0}$, or some equivalent initial data, e.g. $H_{0}$ and $\Omega_{\mathrm{DE}0}$.
@@
In this paper we have followed the majority of work in this area by focussing on generating results from $w_{\mathrm{DE}}$, but there are many other possibilities, as well as further developments and applications, where we describe some in the following list:
@@
  • Instead of using $w_{\mathrm{DE}}$ as the starting point, one can use other variables. For example, as pointed out in footnote 7, our series expansions also give rise to a series expansion in $T$ for $E_{\mathrm{DE}}^{2}$ and thereby also for $E_{\mathrm{DE}}$. This expansion can subsequently be used to obtain various Padé approximants, which can then serve as the starting point for calculating other quantities. Or one can derive series expansions for one of the distance measures and use this for constructing Padé approximants, from which other variables and observables can be derived.
@@
  • For some purposes there might be of interest to obtain an expression for how the scalar field changes during thawing quintessence from the past freezing value $\phi_{*}$ to some arbitrary $e$-fold time $N$. This can be done in several ways, but constructing the $[1,1]_{\phi^{\prime}}$ Padé approximant on the righ hand side of $\phi^{\prime}$ in (43b), i.e. $\sqrt{3(\gamma-1)}\,T/(1+\sigma T)$ where $\sigma=\frac{4}{5}\left(1+\frac{\eta_{*}}{6}\right)$, and then integrating from $\infty$- ∞ to $N$ yields an excellent approximation [^17] It is, e.g., a better approximation than that of Raveri et al. [55] given in equation (72) in appendix A. given by
    $\phi=\phi_{*}+\sqrt{(\gamma-1)/3}\left(\frac{1}{\sigma}\right)\ln{\left(1+\frac{\sigma\Omega_{\phi 0}}{(1-\gamma\Omega_{\phi 0})e^{-3N}}\right)}.$ (66)
@@
  • In the spirit of CPL, it is possible to improve our quintessence approximations even further by viewing $\gamma$ and $\beta$ as free parameters, not determined by the scalar field potential. This, however, means that one only would do curve fitting, which implies less theoretical content and predictive power.
  • We have here focussed on obtaining thawing and tracking quintessence approximations situated in a broader DE context for spatially homogeneous, isotropic and flat spacetimes. A natural next step is to use these approximations in order to interpret and assess the rapidly growing set of increasingly precise observational data. In particular, the present results, or variations thereof, can serve as input in cosmological perturbation theory to bring the results closer to a wide set of observations.
@@ A.1 Thawing quintessence
where inserting (70b) yields $\phi^{\prime}(N)$. This expression agrees with eq. (15) in Agrawal et al. [56] for the exponential potential where $\lambda$ is a constant. Integration, with the condition that ${\phi}={\phi}_{*}$ when $N\rightarrow-\infty$ and $\Omega_{\varphi}\rightarrow 0$ (using $\phi^{\prime}=\Omega_{\phi}^{\prime}d\phi/d\Omega_{\phi}$ and $\Omega_{\phi}^{\prime}\approx 3\Omega_{\phi}(1-\Omega_{\phi})$), results in
@@
Other types of thawing quintessence approximations in the literature are based on series expansions of $V(\phi)$ at some $\phi_{*}$, and thus also of $\lambda(\phi)$ when the latter is regular. An example of this were given by Cahn et al. (2008) [57] who gave approximate expressions for $\phi^{\prime}$ (replacing $\dot{\phi}$ with $\phi^{\prime}$ in their eq. (15)) and $\phi$, although note that $\phi_{*}$ is missing in their eq. (16), thus hiding that you typically are interested in a domain of $\phi_{*}$. These expressions correspond to the present second order expressions in eqs. (43a) and (43b), whose accuracy we subsequently improved by taking Padé approximants.
@@
$\phi$ $=\phi_{*}+\frac{2\lambda_{*}}{3}\left(\frac{(1-\omega^{2})^{\frac{1}{2}(1-K)}\left[(1+\omega)^{K}-(1-\omega)^{K}\right]-2K\omega}{K\omega(K^{2}-1)}\right)$ (73a)
$\phi^{\prime}$ $=\lambda_{*}\frac{(1-\omega^{2})^{\frac{1}{2}(1-K)}\left[(1+K\omega)(1-\omega)^{K}-(1-K\omega)(1+\omega)^{K}\right]}{K\omega(K^{2}-1)}$ (73b)
$1+w_{\phi}$ $=\frac{\lambda^{2}_{*}}{3}(1-\omega^{2})^{1-K}\left[\frac{(1+K\omega)(1-\omega)^{K}-(1-K\omega)(1+\omega)^{K}}{K\omega^{2}(K^{2}-1)}\right]^{2}$ (73c)
@@
In addition, Chiba [53] discussed the approximation given by Crittenden et al. (2007) [59]. Examining the approximations for a variety of potentials we do not find that $K(\phi_{*})\neq 1$ gives better approximations than the simpler $K=1$ expressions, which presumably is due to error cancellations and accumulations in a rather uncontrolled manner during the series of ad hoc approximations that resulted in the $K\neq 1$ expressions.

### v3

@@ 5 Numerical comparisons
Finally, note that changing $\phi_{*}$ so that $\lambda_{*}=\lambda(\phi_{*})$ decreases also decreases the errors, which is due to slower rolling of the scalar field. Note also that the errors of all approximations decrease as $z$ increases, which should not come as a surprise since increasing $z$ entails increased matter domination. Investigations of several other scalar field potentials yield similar results.
(a) Relative errors for $w_{\phi}$
(b) Relative errors for $\Omega_{\mathrm{m}}$
(c) Relative errors for $E_{\phi}$
Figure 1: Thawing quintessence: Relative errors $\left|\Delta F/F\right|$ for $w_{\phi}(z)$, $\Omega_{\mathrm{m}}(z)$, $E_{\phi}(z)$. The dotted curves depict the relative errors of the new $w_{\phi}\approx[0/1]_{w_{\mathrm{DE}}}(T(z))$ based approximations, i.e., (47) with $\beta=\gamma-1$, while the thick dotted curves represent the relative errors of the $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T(z))$ based approximations (47). The dash-dotted curves correspond to the CPL based approximations with the numerically determined values $(w_{0},w_{a})=(-0.9186,-0.0871)$. In Figure (a) the dashed curve is the approximation by Scherrer and Sen [55] given in eq. (88), while the space-dashed curve is the approximation obtained by Chiba [56], see eq. (94). In Fig. (b) and (c) the dashed curves correspond to the $\Lambda$CDM solution given in (19), which is the basis for the approximations of both Scherrer and Sen [55] and Chiba [56].
Figure 2 contains the differences $\Delta(w_{\phi}(z))$, $\Delta(\Omega_{\mathrm{m}}(z))$ and $\Delta(E_{\phi}(z))$ between relative CPL errors and the relative difference between CPL and our approximations, as described in eq. (73). The simple $w_{\phi}\approx[0/1]_{w_{\mathrm{DE}}}(T(z))$ approximant based approximations yields order of estimate descriptions of the CPL approximations but is not sufficiently good to make numerical calculations unnecessary. On the other hand, our analytic $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T(z))$ approximant based approximations are virtually indistinguishable from numerical results, again illustrating that this approximation is so good so that it makes numerical calculations irrelevant.
(a) Difference between relative CPL errors and the relative difference between CPL and our approximations for $w_{\phi}$
(b) Difference between relative CPL errors and the relative difference between CPL and our approximationss for $\Omega_{\mathrm{m}}$
(c) Difference between relative CPL errors and the relative difference between CPL and our approximations for $E_{\phi}$
Figure 2: Thawing quintessence: Difference between relative CPL errors and the relative difference between CPL and our approximations based on $w_{\phi}\approx[0/1]_{w_{\mathrm{DE}}}(T(z))$ (dotted curves) and $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T(z))$ (thick dotted curves), as described in equations (73) for $\Delta(w_{\phi}(z))$, $\Delta(\Omega_{\mathrm{m}}(z))$, $\Delta(E_{\phi}(z))$, emphasizing that the analytic $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T(z))$ based approximations are observationally indistinguishable from numerically based CPL approximations, making such numerical explorations unnecessary.
We finally note that our numerical calculations yield $(w_{0},w_{a})\approx(-0.9186,-0.0871)$, and hence $w^{\mathrm{CPL}}_{\infty}\approx-1.0057$ and $\frac{w_{a}}{w_{0}+1}\approx-1.0700$. Using our $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T(z))$ approximant, which yields equation (69), where we recall that $w_{\infty}=-1$ for thawing quintessence, results in $(w_{0},w_{a})\approx(-0.9183,-0.0791)$ and hence $w_{\infty}^{\mathrm{CPL}}\approx-0.9974$ and $\frac{w_{a}}{w_{0}+1}\approx-0.9682$, which illustrates how good this approximation is.
Tracking quintessence
Thus the derivatives of $\Gamma$ are zero and hence $\Gamma^{(1)}=0$ and $\Gamma^{(2)}=0$ in (57), while, as for all the present tracking quintessence models,
$w_{\infty}=-\frac{2}{2+p},$ (81)
where we will use $w_{\infty}$ rather than $p=-2(1+w_{\infty})/w_{\infty}$ since this leads to somewhat more succinct expressions. The expressions for $\gamma$ and $\beta$ for this case follows straightforwardly from (57) and can be written as
$\gamma$ $=1-\frac{1-w_{\infty}^{2}}{4w_{\infty}^{2}-2w_{\infty}+1},$ (82a)
$\beta$ $=\frac{36w_{\infty}^{4}-21w_{\infty}^{3}+7w_{\infty}^{2}-1}{(4w_{\infty}^{2}-2w_{\infty}+1)(12w_{\infty}^{2}-3w_{\infty}+1)}.$ (82b)
which for $p=1$ yields $w_{\infty}=-2/3$ and $(\gamma,\beta)\approx(0.86486,0.45081)$, while, e.g., $p=0.1$ results in $w_{\infty}=-20/21\approx-0.95238$ and $(\gamma,\beta)\approx(0.98577,0.55145)$. Inserting these results into (58) yields our tracking quintessence approximations (as before, setting first $\beta=\gamma-1$ yields our $w_{\phi}\approx[0/1]_{w_{\mathrm{DE}}}(T(z))$ based approximations). To make numerical comparisons with our approximations we choose the historically important case $p=1$ described above, though this value might be too large to constitute an observationally viable model.
For numerical calculations we use the system (52) and the value $\Omega_{\phi 0}=0.68$, which form a surface in the $(\tilde{\phi},u,v)$ state space that intersects the tracker orbit at a point found numerically, thus identifying where $N=0$, and where the numerical calculations also give us $w_{\phi}=u^{2}-1$, $\Omega_{\mathrm{m}}=1-\Omega_{\phi}=1-3\tilde{\phi}v^{2}$, while $E_{\phi}$ is found numerically by integrating [^16] The present model, like all scalar field models, also admits a 1-parameter set of thawing quintessence attractor solutions in the state space, see [37], which we for brevity do not discuss.
[…]
Finally, note that decreasing $p$ to obtain observationally viable models yields even better results since the maximal errors for the $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T(z))$ based approximations are roughly proportional to $p$, e.g. $p=0.1$ yields a maximal relative error $\sim 0.01\%$ for $E_{\phi}$. Thus, again the $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T)$ based approximations are observationally indistinguishable from numerical results.
@@
Figure 3: Tracking quintessence: Relative errors $\left|\Delta F/F\right|$ for $w_{\phi}(z)$, $\Omega_{\mathrm{m}}(z)=1-\Omega_{\phi}(z)$, $E_{\phi}(z)$. The dotted curves depict the relative errors for the $w_{\phi}\approx[0/1]_{w_{\mathrm{DE}}}(T(z))$ based approximations following from (58), while the thick dotted curves represent the relative errors for the $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T(z))$ based approximations in (58). The dashed-dotted curves correspond to the CPL based approximations with, at the present time, numerically calculated values $(w_{0},w_{a})=(-0.7581,0.1080)$. In Figure (a) the space-dashed and dashed curves describe the Chiba approximation [57], given in (97), truncated at 1st and 2nd order in $\Omega_{\phi}$, respectively. In Fig. (b) and (c) the dashed curves depict the $w$CDM solution given in (19), which form the basis for the Chiba approximations.
We then note that for $p=1$, our numerical calculations yield $(w_{0},w_{a})\approx(-0.7581,0.1080)$, and hence $w^{\mathrm{CPL}}_{\infty}\approx-0.6501$ and $\frac{w_{a}}{w_{0}-w_{\infty}}\approx-1.1812$, where we recall $w_{\infty}=-2/3\approx-0.6667$. The $w_{\phi}\approx[1,1]_{w_{\mathrm{DE}}}$ approximation yields $(w_{0},w_{a})\approx(-0.7519,0.0978)$, and hence $\frac{w_{a}}{w_{0}-w_{\infty}}\approx-1.1474$. Decreasing $p$ yields even better results.
Figure 4 contains the differences $\Delta(w_{\phi}(z))$, $\Delta(\Omega_{\mathrm{m}}(z))$ and $\Delta(E_{\phi}(z))$ between relative CPL errors and the relative difference between CPL and our approximations, as described in eq. (73). Again, the simple $w_{\phi}\approx[0/1]_{w_{\mathrm{DE}}}(T(z))$ approximant based approximations yields order of estimate descriptions of the CPL approximations but not good enough for a detailed description; on the other hand, our analytic $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T(z))$ approximant based approximations are very close to the numerical results, again illustrating that our best approximations are so good so that they make numerical calculations irrelevant.
(a) Absolute difference of relative errors of $w_{\phi}$ for the CPL parametrisation
(b) Absolute difference of relative errors of $\Omega_{\mathrm{m}}$ for the CPL parametrisation
(c) Absolute difference of relative errors of $E_{\phi}$ for the CPL parametrisation
Figure 4: Tracking quintessence: Absolute difference between relative errors, as described in equations (73), for the CPL parametrisation relative to the numerical solution and the $w_{\phi}\approx[0/1]_{w_{\mathrm{DE}}}(T(z))$ (dotted curves) and $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}$ (thick dotted curves) based approximations for $w_{\phi}(z)$, $\Omega_{\mathrm{m}}(z)$, $E_{\phi}(z)$, emphasizing that the analytic $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T(z))$ based approximations are observationally indistinguishable from numerical calculations, making the latter unnecessary.
@@ 6 Discussion
In this paper we have derived simple and accurate approximations for thawing and tracking quintessence. This was accomplished by identifying that thawing and tracking quintessence are associated with ‘attractor solutions’ for which $\lim_{N\rightarrow-\infty}\Omega_{\phi}=0$ (and hence $\lim_{N\rightarrow-\infty}\Omega_{\mathrm{m}}=1$) and $\lim_{N\rightarrow-\infty}w_{\phi}=w_{\infty}$, where $w_{\infty}=-1$ for thawing quintessence while $-1<w_{\infty}<0$ for tracking quintessence. These asymptotic properties served as motivation for situating the derivation of the quintessence approximations in a unifying DE context. It was shown that the conditions $\lim_{N\rightarrow-\infty}\Omega_{\phi}=0$ and $\lim_{N\rightarrow-\infty}w_{\phi}=w_{\infty}\in[-1,0)$ naturally resulted in series expansions in $T=T_{0}\exp(-3w_{\infty}N)$. However, since our goal was to obtain accurate approximations during the whole time period between matter dominance, i.e. when $\Omega_{\phi}\approx 0$ and $\Omega_{\mathrm{m}}\approx 1$, and the present time for thawing and tracking quintessence we used Padé approximants where the $[1/1]_{\Omega_{\mathrm{DE}}}$ Padé approximant for $\Omega_{\mathrm{DE}}$ was used to relate the approximations to the present time initial value $\Omega_{\mathrm{DE}0}$, or, equivalently, $\Omega_{\mathrm{m}0}=\rho_{\mathrm{m}0}/3H_{0}^{2}$.
After the unifying DE section we specialized to thawing and tracking quintessence by deriving two dynamical systems adapted to the two different types of quintessence. This allowed us to insert the DE series expansions into these two dynamical systems and solve for $\gamma$ and $\beta$, which then via the $[0/1]_{w_{\mathrm{DE}}}(T(z))$ and $[1/1]_{w_{\mathrm{DE}}}(T(z))$ based DE expressions, and changing $\Omega_{\mathrm{DE}0}$ to $\Omega_{\phi 0}$, yield corresponding quintessence approximations for $\Omega_{\phi}(z)$ and $E(z)=H(z)/H_{0}$. Note that for a given potential, the parameters in our approximations precisely correspond to the dimensionless parameters in the potential, since they are easily analytically computed from the potential by taking scalar field derivatives; thus, no extra parameters occur in our approximate expressions, although note that thawing quintessence, which exists for all scalar field potentials, also involves a 1-parameter set of solutions described by the freezing value $\phi_{*}$ during the matter dominated epoch. We also showed numerically that the $w_{\phi}(z)\approx[1/1]_{w_{\mathrm{DE}}}(T(z))$ based approximations are an order of magnitude more accurate than any previous analytic approximations in the literature, making numerical calculations for thawing and tracking quintessence unnecessary. Moreover, we showed how one can analytically compute the CPL parameters, again with an accuracy that eliminates the need for numerical calculations of these parameters.
@@
In this paper we have followed the majority of work in this area by focussing on generating results from $w_{\phi}$, but there are many other possibilities, as well as further developments and applications, where we describe some in the following list:
@@
  • Instead of using $w_{\phi}$ as the starting point, one can use other variables. For example, as pointed out in footnote 7, our series expansions also give rise to a series expansion in $T$ for $E_{\mathrm{DE}}^{2}$ and thereby also for $E_{\mathrm{DE}}$ and $E_{\phi}$. This expansion can subsequently be used to obtain various Padé approximants, which can then serve as the starting point for calculating other quantities. Or one can derive series expansions for one of the distance measures and use this for constructing Padé approximants, from which other variables and observables can be derived.
@@
  • For some purposes there might be of interest to obtain an expression for how the scalar field changes during thawing quintessence from the past freezing value $\phi_{*}$ during the matter dominant epoch to some arbitrary $e$-fold time $N$. This can be done in several ways, but constructing the $[1,1]_{\phi^{\prime}}$ Padé approximant on the right hand side of $\phi^{\prime}$ in (41b), i.e. $\sqrt{3(\gamma-1)}\,T/(1+\sigma T)$ where $\sigma=\frac{4}{5}\left(1+\frac{\eta_{*}}{6}\right)$, and then integrating from $\infty$- ∞ to $N$ yields an excellent approximation [^17] It is, e.g., a better approximation than that of Raveri et al. [58] given in equation (90) in appendix A. given by
    $\phi=\phi_{*}+\sqrt{{\textstyle{1\over 3}}(\gamma-1)}\,\left(\frac{1}{\sigma}\right)\ln{\left(1+\frac{\sigma\Omega_{\phi 0}}{(1-\gamma\Omega_{\phi 0})e^{-3N}}\right)}.$ (84)
@@
  • In the spirit of CPL, it is possible to improve our quintessence approximations even further by viewing $\gamma$ and $\beta$ as numerically determined parameters. This, however, means that one (similarly to CPL) would only do curve fitting, which implies less theoretical content and predictive power.
  • We have here focussed on obtaining thawing and tracking quintessence approximations situated in a broader unifying DE context for spatially homogeneous, isotropic and flat spacetimes. A natural next step is to use these approximations in order to interpret and assess the rapidly growing set of increasingly precise observational data. In particular, the present results, or variations thereof, can serve as input in cosmological perturbation theory to bring the results closer to a wide set of observations.
@@ A.1 Thawing quintessence
where inserting (88b) yields $\phi^{\prime}(N)$. This expression agrees with eq. (15) in Agrawal et al. [59] for the exponential potential for which $\lambda$ is a constant. Integration, with the condition that ${\phi}={\phi}_{*}$ when $N\rightarrow-\infty$ and $\Omega_{\varphi}\rightarrow 0$ (using $\phi^{\prime}=\Omega_{\phi}^{\prime}d\phi/d\Omega_{\phi}$ and $\Omega_{\phi}^{\prime}\approx 3\Omega_{\phi}(1-\Omega_{\phi})$), results in
@@
Other types of thawing quintessence approximations in the literature are based on series expansions of $V(\phi)$ at some $\phi_{*}$, and thus also of $\lambda(\phi)$ when the latter is regular at $\phi_{*}$. An example of this were given by Cahn et al. (2008) [60] who gave approximate expressions for $\phi^{\prime}$ (replacing $\dot{\phi}$ with $\phi^{\prime}$ in their eq. (15)) and $\phi$, although note that $\phi_{*}$ is missing in their eq. (16), thus hiding that you typically are interested in a domain of $\phi_{*}$. These expressions correspond to the present second order expressions in eqs. (41a) and (41b), whose accuracy we subsequently improved by taking Padé approximants.
@@
$\phi$ $=\phi_{*}+\frac{2\lambda_{*}}{3}\left(\frac{(1-\omega^{2})^{\frac{1}{2}(1-K)}\left[(1+\omega)^{K}-(1-\omega)^{K}\right]-2K\omega}{K\omega(K^{2}-1)}\right),$ (91a)
$\phi^{\prime}$ $=\lambda_{*}\frac{(1-\omega^{2})^{\frac{1}{2}(1-K)}\left[(1+K\omega)(1-\omega)^{K}-(1-K\omega)(1+\omega)^{K}\right]}{K\omega(K^{2}-1)},$ (91b)
$1+w_{\phi}$ $=\frac{\lambda^{2}_{*}}{3}(1-\omega^{2})^{1-K}\left[\frac{(1+K\omega)(1-\omega)^{K}-(1-K\omega)(1+\omega)^{K}}{K\omega^{2}(K^{2}-1)}\right]^{2},$ (91c)
@@
In addition, Chiba [56] discussed the approximation given by Crittenden et al. (2007) [62].